OpenStudy (anonymous):

(Newton's Law of Motion) A parachutist whose total mass is 100 kg is falling at 50 m/s when her parachute opens. Her speed drops to 7m/s in a vertical distance of 40 m. What total force did her harness have to withstand?

OpenStudy (ash2326):

v^2-u^2=2as here v final velocity is 7m/s u initial velocity is 50 m/s s=40m 7^2-50^2=2a*40 49-2500=80a 80a=-2451 a=-30.6375m/s^2 (-ve because opposite to the motion) Force on the harness=ma =100*30.6375 3063.75 Newtons